# An Upper Bound on the Error Probability of RPA Decoding of Reed-Muller Codes Over the BSC

### v1

@@ IV Analysis on Second-Order RM Codes
In this section, we investigate the performance of the RPA decoder on second-order RM codes, i.e., codes of the form RM$(m,2)$, $m\geq 2$, which are used for transmission over a BSC$(p)$. In particular, our interest is in establishing bounds on the probability that the (iterative) RPA decoder converges to the true transmitted codeword, which is assumed to be drawn uniformly from a second-order RM code.
Let $\mathbf{Y}\in\{0,1\}^{N}$ denote the (random) vector received by passing the input codeword $\mathbf{0}\in\text{RM}(m,2)$ through the BSC$(p)$. We hence have that $\mathbf{Y}\sim\text{Ber}^{\otimes{N}}(p)$. Further, by the definition of the (random) vectors $\mathbf{Y}_{/\mathbb{B}_{i}}$, for any $i\in[N-1]$, it follows that $\mathbf{Y}_{/\mathbb{B}_{i}}\sim\text{Ber}^{\otimes(N/2)}(2p(1-p))$.
@@ IV-A The Performance of the FHT Decoder
Recall from Section II-C that the FHT-based decoder for first-order RM codes is simply an efficient implementation of ML decoding. In this section, we hence directly analyze the behaviour of the ML decoder on first-order RM codes, via a slightly different perspective that is closely related to, but different in form from, the FHT-based decoder.
Now, fix a one-dimensional subspace $\mathbb{B}_{i}$, for some $i\in[N-1]$, as in Algorithm 1; let $\mathbf{b}_{i}\in\{0,1\}^{N}$ denote the non-zero vector in $\mathbb{B}_{i}$. Given the received vector $\mathbf{Y}$, we have that $\mathbf{Y}_{/\mathbb{B}_{i}}\in\text{RM}(m-1,1)$. Consider the ML (or FHT-based) decoder in Step 8 of Algorithm 1 that acts on $\mathbf{Y}_{/\mathbb{B}_{i}}$. First, we transform $\mathbf{Y}_{/\mathbb{B}_{i}}\in\{0,1\}^{N/2}$ to the vector $\mathbf{Y}^{\pm}_{/\mathbb{B}_{i}}\in\{-1,1\}^{N/2}$, via the standard mapping in Section II-A.
@@
$\langle f,g\rangle=1-2\text{dist}(f,g).$
@@
The following key theorem shows that with overwhelming probability, for sufficiently large $N$, the ML decoder, when acting on $\mathbf{Y}^{\pm}_{/\mathbb{B}_{i}}$, produces the correct input codeword $\mathbf{c}\in\text{RM}(m-1,1)$, which is the all-zeros codeword. Let $\eta=\eta(p):=\frac{1}{2}\cdot(1-4p(1-p))$.
@@
$\langle\mathbf{Y}^{\pm}_{/\mathbb{B}_{i}},\chi_{\mathbf{s}}\rangle$ $=\frac{2}{N}\cdot\sum_{\mathbf{x}\in\{0,1\}^{m-1}}\left({Y}^{\pm}_{/\mathbb{B}_{i}}\right)_{\mathbf{x}}$
@@
$\Pr\left[\lvert\langle\mathbf{Y}^{\pm}_{/\mathbb{B}_{i}},\chi_{\mathbf{s}}\rangle\rvert\leq\epsilon\right]$
$=\Pr[|\alpha_{1}-\alpha_{2}|\leq\epsilon]$
@@
$\Pr\left[\mathsf{ML}(\mathbf{Y}^{\pm}_{/\mathbb{B}_{i}})=(\mathbf{0},1),\text{for all }i\in[N-1]\right]$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \geq 1-8N(N-1)\cdot e^{-\frac{N\epsilon^{2}}{8}}.$
@@
$\Pr\left[\widehat{Y}_{/\mathbb{B}_{i}}=\mathbf{0},\text{for all }i\in[N-1]\right]\geq 1-8N(N-1)\cdot e^{-\frac{N\epsilon^{2}}{8}}.$
@@ IV-B The Aggregation Step
Let $\mathcal{G}$ denote the event $\left\{\widehat{Y}_{/\mathbb{B}_{i}}=\mathbf{0},\text{for all }i\in[N-1]\right\}$. Proposition IV.1 shows that $\Pr[\mathcal{G}]\geq 1-8N(N-1)\cdot e^{-\frac{N\epsilon^{2}}{8}}.$ In this subsection, we analyze the performance of the aggregation step, conditioned on the event $\mathcal{G}$. Observe that, conditioned on $\mathcal{G}$, the aggregation step computes, for each $\mathbf{x}\in\{0,1\}^{m}$, the function $\phi(\mathbf{x})=:\phi^{(N)}(\mathbf{x})=\sum_{i=1}^{N-1}\mathds{1}\{Y_{\mathbb{B}_{i}}([\mathbf{x}+\mathbb{B}_{i}])\neq\widehat{Y}_{\mathbb{B}_{i}}([\mathbf{x}+\mathbb{B}_{i}])\}$. Upon conditioning on $\mathcal{G}$, it can be checked that
@@
$\Pr\left[\mathsf{Flip}^{(N)}(\mathbf{x})=\mathds{1}\left\{\overline{\phi}_{\infty}(\mathbf{x})>\frac{1}{2}\right\}\bigg{|}\mathcal{G}\right]\geq$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ 1-16N(N-1)\cdot e^{-\frac{N\epsilon^{2}}{8}}.$
@@
Now, via Hoeffding’s inequality [31] (see also [32, Thm. 2.2.6]), with probability at least $1-2\cdot e^{-2(N-1)\epsilon^{2}}$, we have
$\left\lvert\frac{1}{N-1}\cdot\sum_{\mathbf{z}\neq\mathbf{x}}Y_{\mathbf{z}}-p\right\rvert\leq\epsilon.$
Substituting this back in (5), we obtain that with probability at least $1-2\cdot e^{-2(N-1)\epsilon^{2}}$, we must have $\left\lvert\overline{\phi}^{(N)}(\mathbf{x})-\overline{\phi}_{\infty}(\mathbf{x})\right\rvert\leq\epsilon,$ thereby proving the lemma. ∎
@@
$\stackrel{{\scriptstyle(a)}}{{\geq}}\Pr\left[\left\lvert\mathds{1}\left\{\overline{\phi}(\mathbf{x})>\frac{1}{2}\right\}-\mathds{1}\left\{\overline{\phi}_{\infty}(\mathbf{x})>\frac{1}{2}\right\}\right\rvert\leq\epsilon\ \bigg{|}\ \overline{\mathcal{G}}\right]\cdot\Pr[\overline{\mathcal{G}}]$
@@
$\geq 1-\delta-\delta^{\prime}\geq 1-2\delta^{\prime}.$
Here, we define $\mathcal{G}^{c}:=\left\{\widehat{Y}_{/\mathbb{B}_{i}}\neq\mathbf{0},\text{for some }i\in[N-1]\right\}$, which leads to the first inequality above by simple manipulations of the law of total probability. The second inequality above then makes use of Proposition IV.1.
@@
$\Pr\left[\mathsf{Flip}^{(N)}(\mathbf{x})=\mathds{1}\left\{\overline{\phi}_{\infty}(\mathbf{x})>\frac{1}{2}\right\},\text{ for all $\mathbf{x}$}\ \bigg{|}\ \mathcal{G}\right]$
$\geq 1-16N^{2}(N-1)\cdot e^{-\frac{N\epsilon^{2}}{8}}.$ (6)
@@
$\Pr\left[\mathsf{Flip}^{(N)}=\mathbf{Y}\right]$ (7)
$\geq\Pr\left[\mathsf{Flip}^{(N)}=\mathbf{Y}\ \bigg{|}\ \mathcal{G}\right]\cdot\Pr[\mathcal{G}]$ (8)
@@
Proof.
From Proposition IV.2, for any $\mathbf{x}\in\{0,1\}^{m}$, we have that $\Pr\left[\mathsf{Flip}^{(N)}(\mathbf{x})=Y_{\mathbf{x}}\ \bigg{|}\ \mathcal{G}\right]\geq 1-16N(N-1)\cdot e^{-\frac{N\epsilon^{2}}{8}}.$ Via Markov’s inequality, this then allows us to obtain an upper bound on the probability that, conditioned on the event $\mathcal{G}$, the vector $\textsf{Flip}^{(N)}$ differs from $\mathbf{Y}$ in at least $2^{m-3}$ positions, where $22^{m-3}=N/8=d_{\text{min}}(\text{RM}(m,2))/$ 2. Indeed, we have that
@@
By calculations similar to those in (8)–(9), we obtain $\Pr\left[w\left(\textsf{Flip}^{(N)}\oplus\mathbf{Y}\right)\geq N/8\right]\geq 1-256N^{2}e^{-\frac{N\epsilon^{2}}{8}}$. Now, note that if $\textsf{Flip}^{(N)}$ and $\mathbf{Y}$ differ in at most $N/8-1$ positions, we have that at the end of one RPA iteration, $w\left(\overline{\mathbf{Y}}\right)\leq d_{\text{min}}(\text{RM}(m,2))/2-1$. This hence implies, via arguments as in [25, Lemma 1] that the RPA decoder, in the next iteration, can correct all errors present in $\overline{\mathbf{Y}}$, resulting in convergence to the correct all-zeros codeword, at the end of two iterations. ∎
@@ V Extension to Higher-Order RM Codes
$\mathbf{Y}_{/\mathbb{S}^{(1)}}:=\left(\mathbf{Y}_{/\mathbb{S}^{(1)}}(T):\ T\in\{0,1\}^{m}/\mathbb{S}^{(1)}\right).$
@@
$(\displaystyle\mathbf{Y}_{/\left(\mathbb{S}^{(1)},\ldots,\mathbb{S}^{(j+1)}\right)}:=\bigg{(}$$\left(\mathbf{Y}_{/\left(\mathbb{S}^{(1)},\mathbb{S}^{(2)},\ldots,\mathbb{S}^{(j)}\right)}\right)_{/\mathbb{S}^{(j+1)}}(T):$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ T\in\{0,1\}^{m-j}/\mathbb{S}^{(j+1)}\bigg{)}.$ (10)
For any (random) codeword $\mathbf{C}\in\text{RM}(m,r)$, one can define projections $\mathbf{C}_{/\left(\mathbb{S}^{(1)},\mathbb{S}^{(2)},\ldots,\mathbb{S}^{(j)}\right)}$, for $1\leq j\leq m,$ analogously; via arguments made in [22, Lemma 1], it can be shown that $\mathbf{C}_{/\left(\mathbb{S}^{(1)},\mathbb{S}^{(2)},\ldots,\mathbb{S}^{(j)}\right)}\in\text{RM}(m-j,r-j)$, if $\mathbf{C}\in\text{RM}(m,r)$.
The definitions above, along with the recursive structure of the RPA decoder (see Algorithm 1) motivate the definition of a “projection-aggregation tree"—a simple, visual device that helps us recursively compute the probabilities of error in decoding for higher-order RM codes. Formally, a projection-aggregation tree is a rooted tree, with the root (at level 00) being the code RM$(m,r)$. For every level $30\leq i\leq r-$ 3, each node at level $i$ has $\frac{N}{2^{i}}-1$ children at level $i+1$, each of which is an RM$(m-i-1,r-i-1)$ code; here, $N:=2^{m}$. Clearly, the leaves of the projection-aggregation tree are RM$(m-r+2,2)$ codes. Figure 1 shows a pictorial depiction of a projection-aggregation tree. Each path in the tree from the root to a leaf node corresponds to a fixed sequence $\mathbb{S}_{1},\ldots,\mathbb{S}_{r-2}$ of one-dimensional subspaces as above. “Moving down" from a parent node to a child node in this tree signifies a projection step and “moving up" from a child node to its parent signifies an aggregation step.
Now, for any $j\in[r-2]$, let $p^{(j)}:=\frac{1}{2}\cdot(1-(1-2p)^{2^{j}})$ and $N^{(j)}:=N/2^{j}$. Since $\mathbf{Y}\sim\text{Ber}^{\otimes N}(p)$, the following simple lemma holds:
@@
First, observe that the vector $\mathbf{Y}_{/\left(\mathbb{S}^{(1)},\ldots,\mathbb{S}^{(j)}\right)}$ is of length $N/2^{j}=N^{(j)}$, by construction. Further, by the disjointness of distinct cosets, we obtain from (V) that each coordinate of $\mathbf{Y}_{/\left(\mathbb{S}^{(1)},\ldots,\mathbb{S}^{(j)}\right)}$ is the sum, modulo $2$, of $2^{j}$ distinct coordinates of $\mathbf{Y}$. This implies that any coordinate of $\mathbf{Y}_{/\left(\mathbb{S}^{(1)},\ldots,\mathbb{S}^{(j)}\right)}$ takes the value $1$ with probability equal to the probability that the parity, or the sum modulo $2$ of $2^{j}$ Ber$(p)$ random variables, equals $1$, which in turn equals $p^{(j)}$. Moreover, distinct coordinates of $\mathbf{Y}_{/\left(\mathbb{S}^{(1)},\ldots,\mathbb{S}^{(j)}\right)}$ correspond to modulo $2$ sums of disjoint collections of coordinates in $\mathbf{Y}$, thereby yielding the statement of the lemma. ∎
@@
$\left(\mathbf{Y}_{/\left(\overline{\mathbb{S}}^{(1)},\ldots,\overline{\mathbb{S}}^{(r-2)},\mathbb{S}^{(r-1)}\right)}:\ \mathbb{S}^{(r-1)}\in\{0,1\}^{m-r+2}\right)$
@@
$\Pr$ $\left[\textsc{Agg}^{(r-2)}\left(\overline{\mathbb{S}}^{(1)},\ldots,\overline{\mathbb{S}}^{(r-2)}\right)=\mathbf{0}\right]\geq$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ 1-32\left(N^{(r-2)}\right)^{3}\cdot e^{-\frac{N^{(r-2)}\epsilon^{2}}{8}}.$
@@
$\Pr$ $\left[\widehat{\mathbf{Y}}_{/\left({\overline{\mathbb{S}}^{(1)},\ldots,\overline{\mathbb{S}}^{(r-3)},\mathbb{S}^{(r-2)}}\right)}=\mathbf{0},\ \text{for all ${\mathbb{S}}^{(r-2)}$}\right]\geq$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ 1-32\left(N^{(r-2)}\right)^{4}\cdot e^{-\frac{N^{(r-2)}\epsilon^{2}}{8}},$ (11)
@@
$\Pr\left[\widehat{\mathbf{Y}}_{/\left({\overline{\mathbb{S}}^{(1)},\ldots,\overline{\mathbb{S}}^{(i-2)},\mathbb{S}^{(i-1)}}\right)}=\mathbf{0}\right]$
$\geq 1-16\left(N^{(i-1)}\right)^{2}\left(N^{(i-1)}-1\right)\cdot e^{-\frac{N^{(i-1)}\epsilon^{2}}{8}}-$
@@
From Theorem V.1, we see that for any $\epsilon<\eta(p^{(r-2)})$,
@@
$P_{\text{err}}(\text{RM}(m,r))$
$\leq 16N^{2}(N-1)\cdot e^{-\frac{N\epsilon^{2}}{8}}+2N^{r+1}\cdot e^{-\frac{N^{(r-2)}\epsilon^{2}}{8}}$

### v2

@@ IV Analysis on Second-Order RM Codes
In this section, we investigate the performance of the RPA decoder on second-order RM codes, i.e., codes of the form RM$(m,2)$, $m\geq 2$, which are used for transmission over a BSC$(p)$. Such codes form the “base case" in our recursive analysis of the performance of RPA decoding, in that they are the smallest codes that the RPA decoder performs both the FHT decoding and aggregation steps on. Our interest is hence in establishing bounds on the probability that the (iterative) RPA decoder converges to the true transmitted codeword.
Let $\mathbf{Y}\in\{0,1\}^{N}$ denote the (random) vector received by passing the input codeword $\mathbf{0}\in\text{RM}(m,2)$ through the BSC$(p)$ (recall the argument in Section III that shows that we can restrict attention to the all-zeros input codeword). We hence have that $\mathbf{Y}\sim\text{Ber}^{\otimes{N}}(p)$. Further, by the definition of the (random) vectors $\mathbf{Y}_{/\mathbb{B}_{i}}$, for any $i\in[N-1]$, it follows that $\mathbf{Y}_{/\mathbb{B}_{i}}\sim\text{Ber}^{\otimes(N/2)}(2p(1-p))$. We recall here that the number of projections is $n_{1,m}=N-1$.
@@ IV-A The Performance of the FHT Decoder
Recall from Section II-C that the FHT-based decoder for first-order RM codes is simply an efficient implementation of ML decoding. In this section, we hence directly analyze the behaviour of the ML decoder on first-order RM codes, via a slightly different perspective that is closely related to, but different in form from, the FHT-based decoder. We mention that [18] also presents (tight) upper bounds on the error probability of ML decoding of first-order RM codes; however, we adopt a simpler analysis that is based on standard concentration inequalities.
Now, fix a one-dimensional subspace $\mathbb{B}_{i}$, for some $i\in[N-1]$, as in Algorithm 1; let $\mathbf{b}_{i}\in\{0,1\}^{N}$ denote the non-zero vector in $\mathbb{B}_{i}$. Given the received vector $\mathbf{Y}$, we have that $\mathbf{Y}_{/\mathbb{B}_{i}}$ is a noisy version of a codeword in $\text{RM}(m-1,1)$. Consider the ML (or FHT-based) decoder in Step 8 of Algorithm 1 that acts on $\mathbf{Y}_{/\mathbb{B}_{i}}$. First, we transform $\mathbf{Y}_{/\mathbb{B}_{i}}\in\{0,1\}^{N/2}$ to the vector $\mathbf{Y}^{\pm}_{/\mathbb{B}_{i}}\in\{-1,1\}^{N/2}$, via the standard mapping in Section II-A.
@@
$\langle f,g\rangle=1-2\cdot\text{dist}(f,g).$
@@
The following key theorem shows that with very high probability, for sufficiently large $N$, the ML decoder, when acting on $\mathbf{Y}^{\pm}_{/\mathbb{B}_{i}}$, produces the correct input codeword $\mathbf{c}\in\text{RM}(m-1,1)$, which is the all-zeros codeword. Let $\eta=\eta(p):=\frac{1}{2}\cdot(1-4p(1-p))$.
@@
$\langle\mathbf{Y}^{\pm}_{/\mathbb{B}_{i}},\chi_{\mathbf{s}}\rangle$ $=\frac{2}{N}\cdot\sum_{\mathbf{x}\in\{0,1\}^{m-1}}\left({Y}^{\pm}_{/\mathbb{B}_{i}}\right)_{\mathbf{x}}\cdot(-1)^{\mathbf{x}\cdot\mathbf{s}}$
@@
$\Pr\left[\lvert\langle\mathbf{Y}^{\pm}_{/\mathbb{B}_{i}},\chi_{\mathbf{s}}\rangle\rvert\leq\epsilon\right]$ $=\Pr[|\alpha_{1}-\alpha_{2}|\leq\epsilon]$
@@
$\Pr\left[\mathsf{ML}(\mathbf{Y}^{\pm}_{/\mathbb{B}_{i}})=(\mathbf{0},1),\text{for all }i\in[N-1]\right]\geq 1-8N(N-1)\cdot e^{-\frac{N\epsilon^{2}}{8}}.$
@@
$\Pr\left[\mathbf{\widehat{Y}}_{/\mathbb{B}_{i}}=\mathbf{0},\text{for all }i\in[N-1]\right]\geq 1-8N(N-1)\cdot e^{-\frac{N\epsilon^{2}}{8}}.$
@@ IV-B The Aggregation Step
Let $\mathcal{G}$ denote the event $\left\{\mathbf{\widehat{Y}}_{/\mathbb{B}_{i}}=\mathbf{0},\text{for all }i\in[N-1]\right\}$. Proposition IV.1 shows that $\Pr[\mathcal{G}]\geq 1-8N(N-1)\cdot e^{-\frac{N\epsilon^{2}}{8}}.$ In this subsection, we analyze the performance of the aggregation step, conditioned on the event $\mathcal{G}$. Observe that, conditioned on $\mathcal{G}$, the aggregation step computes, for each $\mathbf{x}\in\{0,1\}^{m}$, the function $\phi(\mathbf{x})=:\phi^{(N)}(\mathbf{x})=\sum_{i=1}^{N-1}\mathds{1}\{Y_{/\mathbb{B}_{i}}([\mathbf{x}+\mathbb{B}_{i}])\neq\widehat{Y}_{/\mathbb{B}_{i}}([\mathbf{x}+\mathbb{B}_{i}])\}$. Upon conditioning on $\mathcal{G}$, it can be checked that $\phi(\mathbf{x})$ obeys
@@
$\Pr\left[\mathsf{Flip}^{(N)}(\mathbf{x})=\mathds{1}\left\{\overline{\phi}_{\infty}(\mathbf{x})>\frac{1}{2}\right\}\bigg{|}\ \mathcal{G}\right]\geq 1-16N(N-1)\cdot e^{-\frac{N\epsilon^{2}}{8}}.$
@@
Now, conditioned on any fixed value $y_{\mathbf{x}}\in\{0,1\}$, we have via Hoeffding’s inequality [31] (see also [32, Thm. 2.2.6]), that with probability at least $1-2\cdot e^{-2(N-1)\epsilon^{2}}$,
$\left\lvert\frac{1}{N-1}\cdot\sum_{\mathbf{z}\neq\mathbf{x}}Y_{\mathbf{z}}-p\right\rvert\leq\epsilon,$
or equivalently, that
$\left\lvert\left(1-\frac{1}{N-1}\cdot\sum_{\mathbf{z}\neq\mathbf{x}}Y_{\mathbf{z}}\right)-(1-p)\right\rvert\leq\epsilon.$
We mention that the above result uses the fact that $Y_{\mathbf{x}}$ is independent of $\{Y_{\mathbf{z}}\}_{\mathbf{z}\neq\mathbf{x}}$. Substituting this back in (6) and using the law of total probability, we obtain that with probability at least $1-2\cdot e^{-2(N-1)\epsilon^{2}}$, we must have $\left\lvert\overline{\phi}^{(N)}(\mathbf{x})-\overline{\phi}_{\infty}(\mathbf{x})\right\rvert\leq\epsilon,$ thereby proving the lemma. ∎
@@
$\Pr\left[\left\lvert\mathds{1}\left\{\overline{\phi}(\mathbf{x})>\frac{1}{2}\right\}-\mathds{1}\left\{\overline{\phi}_{\infty}(\mathbf{x})>\frac{1}{2}\right\}\right\rvert\leq\epsilon\right]$ $\stackrel{{\scriptstyle(a)}}{{\geq}}\Pr\left[\left\lvert\mathds{1}\left\{\overline{\phi}(\mathbf{x})>\frac{1}{2}\right\}-\mathds{1}\left\{\overline{\phi}_{\infty}(\mathbf{x})>\frac{1}{2}\right\}\right\rvert\leq\epsilon\ \bigg{|}\ \overline{\mathcal{G}}\right]\cdot\Pr[\overline{\mathcal{G}}]$
@@
$\geq 1-\delta-\delta^{\prime}\geq 1-2\delta^{\prime},$
since $\delta^{\prime}\geq\delta$, for all $N\geq 2$. Here, we define $\mathcal{G}^{c}:=\left\{\widehat{Y}_{/\mathbb{B}_{i}}\neq\mathbf{0},\text{for some }i\in[N-1]\right\}$, which leads to the first inequality above by simple manipulations of the law of total probability. The second inequality above then makes use of Proposition IV.1.
@@
$\Pr\left[\mathsf{Flip}^{(N)}(\mathbf{x})=\mathds{1}\left\{\overline{\phi}_{\infty}(\mathbf{x})>\frac{1}{2}\right\},\text{ for all $\mathbf{x}$}\ \bigg{|}\ \mathcal{G}\right]$ $\geq 1-16N^{2}(N-1)\cdot e^{-\frac{N\epsilon^{2}}{8}}.$ (7)
@@
$\Pr\left[\mathsf{Flip}^{(N)}=\mathbf{Y}\right]$ $\geq\Pr\left[\mathsf{Flip}^{(N)}=\mathbf{Y}\ \bigg{|}\ \mathcal{G}\right]\cdot\Pr[\mathcal{G}]$ (8)
@@
Proof of Lemma IV.5.
From Proposition IV.2, for any $\mathbf{x}\in\{0,1\}^{m}$, we have that $\Pr\left[\mathsf{Flip}^{(N)}(\mathbf{x})=Y_{\mathbf{x}}\ \bigg{|}\ \mathcal{G}\right]\geq 1-16N(N-1)\cdot e^{-\frac{N\epsilon^{2}}{8}}.$ Via Markov’s inequality, this then allows us to obtain an upper bound on the probability that, conditioned on the event $\mathcal{G}$, the vector $\textsf{Flip}^{(N)}$ differs from $\mathbf{Y}$ in at least $2^{m-3}$ positions, where $22^{m-3}=N/8=d/$ 2. Indeed, we have that
@@
By calculations similar to those in (8)–(9), we obtain $\Pr\left[w\left(\textsf{Flip}^{(N)}\oplus\mathbf{Y}\right)\geq N/8\right]\geq 1-256N^{2}e^{-\frac{N\epsilon^{2}}{8}}$. Now, note that if $\textsf{Flip}^{(N)}$ and $\mathbf{Y}$ differ in at most $N/8-1$ positions, we have that at the end of one RPA iteration, $w\left(\overline{\mathbf{Y}}\right)\leq d_{\text{min}}(\text{RM}(m,2))/2-1$. This hence implies, via arguments as in [33, Lemma 1] that the RPA decoder, in the next iteration, can correct all errors present in $\overline{\mathbf{Y}}$, resulting in convergence to the correct all-zeros codeword, at the end of two iterations. ∎
@@ V Extension to Higher-Order RM Codes
$\mathbf{Y}_{/\mathbb{S}^{(1)}}:=\left({Y}_{/\mathbb{S}^{(1)}}(T):\ T\in\{0,1\}^{m}/\mathbb{S}^{(1)}\right).$
@@
$(\displaystyle\mathbf{Y}_{/\left(\mathbb{S}^{(1)},\ldots,\mathbb{S}^{(j+1)}\right)}:=\bigg{(}$$\left({Y}_{/\left(\mathbb{S}^{(1)},\mathbb{S}^{(2)},\ldots,\mathbb{S}^{(j)}\right)}\right)_{/\mathbb{S}^{(j+1)}}(T):T\in\{0,1\}^{m-j}/\mathbb{S}^{(j+1)}\bigg{)}.$ (10)
For any (random) codeword $\mathbf{C}\in\text{RM}(m,r)$, one can define projections $\mathbf{C}_{/\left(\mathbb{S}^{(1)},\mathbb{S}^{(2)},\ldots,\mathbb{S}^{(j)}\right)}$, for $1\leq j\leq m,$ analogously; via arguments made in [23, Lemma 1], it can be shown that $\mathbf{C}_{/\left(\mathbb{S}^{(1)},\mathbb{S}^{(2)},\ldots,\mathbb{S}^{(j)}\right)}\in\text{RM}(m-j,r-j)$, if $\mathbf{C}\in\text{RM}(m,r)$.
The definitions above, along with the recursive structure of the RPA decoder (see Algorithm 1) motivate the definition of a “projection-aggregation tree"—a simple, visual device that helps us recursively compute the probabilities of error in decoding for higher-order RM codes. Formally, a projection-aggregation tree is a rooted tree, with the root (at level 00) being the code RM$(m,r)$. For every level $30\leq i\leq r-$ 3, each node at level $i$ has $\frac{N}{2^{i}}-1$ children at level $i+1$, each of which is an RM$(m-i-1,r-i-1)$ code; here, we recall that $N=2^{m}$. Clearly, the leaves of the projection-aggregation tree are RM$(m-r+2,2)$ codes. Figure 1 shows a pictorial depiction of a projection-aggregation tree. Each path in the tree from the root to a leaf node corresponds to a fixed sequence $\mathbb{S}_{1},\ldots,\mathbb{S}_{r-2}$ of one-dimensional subspaces as above. “Moving down" from a parent node to a child node in this tree signifies a projection step and “moving up" from a child node to its parent signifies an aggregation step.
Now, for any $j\in[r-2]$, recall that $p^{(j)}=\frac{1}{2}\cdot(1-(1-2p)^{2^{j}})$ and let $N^{(j)}:=N/2^{j}$. Since $\mathbf{Y}\sim\text{Ber}^{\otimes N}(p)$, the following simple lemma holds:
@@
First, observe that the vector $\mathbf{Y}_{/\left(\mathbb{S}^{(1)},\ldots,\mathbb{S}^{(j)}\right)}$ is of length $N/2^{j}=N^{(j)}$, by construction. Further, by the disjointness of distinct cosets, we obtain from (10) that each coordinate of $\mathbf{Y}_{/\left(\mathbb{S}^{(1)},\ldots,\mathbb{S}^{(j)}\right)}$ is the sum, modulo $2$, of $2^{j}$ distinct coordinates of $\mathbf{Y}$. This implies that any coordinate of $\mathbf{Y}_{/\left(\mathbb{S}^{(1)},\ldots,\mathbb{S}^{(j)}\right)}$ takes the value $1$ with probability equal to the probability that the parity, or the sum modulo $2$ of $2^{j}$ Ber$(p)$ random variables, equals $1$, which in turn equals $p^{(j)}$. Moreover, distinct coordinates of $\mathbf{Y}_{/\left(\mathbb{S}^{(1)},\ldots,\mathbb{S}^{(j)}\right)}$ correspond to modulo $2$ sums of disjoint collections of coordinates in $\mathbf{Y}$, thereby yielding the statement of the lemma. ∎
@@
$\left(\mathbf{Y}_{/\left(\overline{\mathbb{S}}^{(1)},\ldots,\overline{\mathbb{S}}^{(r-2)},\mathbb{S}^{(r-1)}\right)}:\ \mathbb{S}^{(r-1)}\in\{0,1\}^{m-r+2}\right).$
@@
$\Pr$ $\left[\textsc{Agg}^{(r-2)}\left(\overline{\mathbb{S}}^{(1)},\ldots,\overline{\mathbb{S}}^{(r-2)}\right)=\mathbf{0}\right]\geq 1-32\left(N^{(r-2)}\right)^{3}\cdot e^{-\frac{N^{(r-2)}\epsilon^{2}}{8}}.$
@@
$\Pr$ $\left[\widehat{\mathbf{Y}}_{/\left({\overline{\mathbb{S}}^{(1)},\ldots,\overline{\mathbb{S}}^{(r-3)},\mathbb{S}^{(r-2)}}\right)}=\mathbf{0},\ \text{for all ${\mathbb{S}}^{(r-2)}$}\right]\geq 1-32\left(N^{(r-2)}\right)^{4}\cdot e^{-\frac{N^{(r-2)}\epsilon^{2}}{8}},$ (11)
@@
$\Pr\left[\widehat{\mathbf{Y}}_{/\left({\overline{\mathbb{S}}^{(1)},\ldots,\overline{\mathbb{S}}^{(i-2)},\mathbb{S}^{(i-1)}}\right)}=\mathbf{0}\right]$ $\geq 1-16\left(N^{(i-1)}\right)^{2}\left(N^{(i-1)}-1\right)\cdot e^{-\frac{N^{(i-1)}\epsilon^{2}}{8}}-$
@@
From Theorem V.1, we see that for any $\epsilon<\eta(p^{(r-2)})=\eta(\overline{p})$,
@@
$P_{\text{err}}(\text{RM}(m,r))$ $\leq 16N^{2}(N-1)\cdot e^{-\frac{N\epsilon^{2}}{8}}+2N^{r+1}\cdot e^{-\frac{N^{(r-2)}\epsilon^{2}}{8}}$